OpenStudy (anonymous):

4. If x = 20 when y = -4, find y when x = 10.

OpenStudy (ahsome):

Is this a linear equation, @dianolove?

OpenStudy (ahsome):

Wait, couldn't you do: \[\begin{align} y&=kx\\ -4&=k\times20\\ k&=\dfrac{-1}{5} \end{align}\]

OpenStudy (anonymous):

this is the question it asked me Suppose that x and y vary inversely. Write a function that models each inverse variation. Then use the equation to solve the problem.

OpenStudy (anonymous):

Let's solve your equation step-by-step. 20x−4x=10 Step 1: Simplify both sides of the equation. 20x−4x=10 Simplify: (Show steps) 16x=10 Step 2: Divide both sides by 16. 16x16=1016 x=58 Answer: x=58

OpenStudy (anonymous):

do you under stand

OpenStudy (anonymous):

20x−4y=10

OpenStudy (phi):

*** x and y vary inversely*** means you can write the formula \[ y = \frac{k}{x} \]

OpenStudy (anonymous):

it is a inverse variation and your solving for y not x

OpenStudy (phi):

that formula is useful if we know what "k" is. to find k, we use the info If x = 20 when y = -4, find y when x = 10.

OpenStudy (phi):

in other words, replace x with 20 and y with -4 into the formula \[ -4= \frac{k}{20} \] and "solve for k" by multiplying both sides by 20

OpenStudy (phi):

can you find what k is ?

OpenStudy (anonymous):

k is -80

OpenStudy (phi):

yes, so now you know the formula (for this problem) is \[ y = \frac{-80}{x} \] find y when x = 10. to find y, replace x with 10 in the formula and simplify

OpenStudy (anonymous):

y=-8

OpenStudy (phi):

yes

OpenStudy (anonymous):

thanks phi will you keep helping me with these i got another one

OpenStudy (phi):

ok

OpenStudy (anonymous):

If x = 5 when , find y when x = 10.

OpenStudy (phi):

missing some info

OpenStudy (anonymous):

5. If x = 5 when y=-1 third , find y when x = 10.

OpenStudy (phi):

do you mean if x=5 when y = \( - \frac{1}{3} \) ?

OpenStudy (anonymous):

sorry mines will not do that but yes

OpenStudy (anonymous):

is k -30

OpenStudy (phi):

you could type it as y= -1/3 \[ y = \frac{k}{x} \\ - \frac{1}{3} = \frac{k}{5} \] to find k, multiply both sides by 5 \[ - \frac{1}{3} \cdot 5 = \frac{k}{5}\cdot 5 \\ - \frac{1}{3} \cdot 5 = k \cdot \frac{5}{5} \\ - \frac{5}{3} = k \]

OpenStudy (anonymous):

-15

OpenStudy (phi):

what is -15 ? k is -5/3

OpenStudy (anonymous):

oh i thought you said t multiply -5 into 3 , but what do you do now divide -5/3 into 10

OpenStudy (phi):

you write the formula \[ y = \frac{k}{x} \] and it might be easier to understand if you write it this way \[ y = k \cdot \frac{1}{x} \] which means the same thing. now replace k with -5/3 \[ y = - \frac{5}{3} \cdot \frac{1}{x} \]

OpenStudy (phi):

find y when x = 10. that means put in 10 for x in the formula \[ y = - \frac{5}{3} \cdot \frac{1}{x} \\ y = - \frac{5}{3} \cdot \frac{1}{10}\] and simplify

OpenStudy (anonymous):

y=6

OpenStudy (phi):

if you use a calculator on \[ - \frac{5}{3} \cdot \frac{1}{10} \] what do you get ?

OpenStudy (anonymous):

-1/6

OpenStudy (phi):

yes, y= -1/6 when x=10. It would be good to be able to do that without the calculator. when you multiply fractions, you multiply top times top and bottom times bottom \[ - \frac{5}{3} \cdot \frac{1}{10} = - \frac{5 \cdot 1 }{3 \cdot 10} \] we could multiply the bottom 3*10 to get 30 \[ - \frac{5}{30} \] and then simplify to get -1/6 or, it's easier to notice that 10 is 2*5 \[ - \frac{5 \cdot 1 }{3 \cdot 10} = - \frac{5 \cdot 1 }{3 \cdot 2 \cdot 5} \] and notice we have a 5 up top and down below, and they simplify to 5/5 = 1 \[ - \frac{1 }{3 \cdot 2} = - \frac{1}{6} \]

OpenStudy (anonymous):

what if x=-4/15 when y=-105,find x when y=4

OpenStudy (phi):

\[ y = k \cdot \frac{1}{x} \\ -105= k \cdot \frac{1}{-\frac{4}{15}} \]

OpenStudy (phi):

multiply both sides by -4/15,

OpenStudy (anonymous):

k=28

OpenStudy (phi):

so \[y = \frac{28}{x} \] find x when y=4 replace y with y \[ 4 = \frac{28}{x} \]

OpenStudy (phi):

first multiply both sides by x then divide both sides by 4

OpenStudy (anonymous):

7

OpenStudy (phi):

yes. \[ 4 x= \frac{28}{x} \cdot x \\ 4x = 28 \\\frac{4x}{4}= \frac{28}{4} \\x=7\]

OpenStudy (anonymous):

thanks are you getting tired of me asking questions

OpenStudy (phi):

no, but I do have to go. make a new post and someone will help you.

OpenStudy (anonymous):

ok thanks